### v1

# Data-Driven Global Sensitivity Analysis for Engineering Design Based on Individual Conditional Expectations

Abstract
Explainable machine learning techniques have gained increasing attention in engineering applications, especially in aerospace design and analysis, where understanding how input variables influence data-driven models is essential. Partial Dependence Plots (PDPs) are widely used for interpreting black-box models by showing the average effect of an input variable on the prediction. However, their global sensitivity metric can be misleading when strong interactions are present, as averaging tends to obscure interaction effects. To address this limitation, we propose a global sensitivity metric based on Individual Conditional Expectation (ICE) curves. The method computes the expected feature importance across ICE curves, along with their standard deviation, to more effectively capture the influence of interactions. We provide a mathematical proof demonstrating that the PDP-based sensitivity is a lower bound of the proposed ICE-based metric under truncated orthogonal polynomial expansion. In addition, we introduce an ICE-based correlation value to quantify how interactions modify the relationship between inputs and the output. Comparative evaluations were performed on three cases: a 5-variable analytical function, a 5-variable wind-turbine fatigue problem, and a 9-variable airfoil aerodynamics case, where ICE-based sensitivity was benchmarked against PDP, SHapley Additive exPlanations (SHAP), and Sobol’ indices. The results show that ICE-based feature importance provides richer insights than the traditional PDP-based approach, while visual interpretations from PDP, ICE, and SHAP complement one another by offering multiple perspectives.
keywords:
Global sensitivity analysis, Individual Conditional Expectations, Partial Dependence Plot, Engineering Design, Aerodynamics
journal: Aerospace Science and Technology
1 Introduction
High-fidelity simulations are essential in modern engineering design. Therefore, aerospace engineers often rely on computational fluid dynamics to assess and optimize aerodynamic performance, enabling accurate prediction of behavior under varied conditions and supporting informed, performance- and safety-driven decisions [1]. As engineering problems grow more complex, especially in aerospace, precise simulation data is important. Yet high-fidelity simulations remain challenging for design exploration because engineers can only afford to run a limited number of computationally intensive simulations. To address this, a surrogate model that estimates the relationship between input or design variables with the Quantity of Interest (QoI) plays an important role in accomplishing this task. The surrogate can be used for various purposes such as optimization [2] and uncertainty analysis [3]. Various types of surrogate models exist, among popular ones in engineering, including polynomial regression (including polynomial chaos expansion (PCE) [4, 5]), Kriging/Gaussian process regression [6, 7], and deep neural network [8] to reduce the computational cost of inference.
Sensitivity analysis, particularly Global Sensitivity Analysis (GSA), supports engineering endeavours by identifying which variables are most and least influential [9]. In the context of engineering design exploration, this information is important for setting design priorities, such as determining which variables should be prioritized. For instance, in the field of aerodynamics, GSA aids in analyzing the impact of geometric or kinematic parameters on the aerodynamic performance of a multi-element airfoil [10], a hovering flapping wing [11], and a high-pressure capturing wing [12]. Common techniques in engineering include variance-based decomposition methods, such as Sobol' indices [13], which measure the individual effects of input variables and their interactions on the QoI. Although variance-based decomposition is valuable, it only quantifies how much each input contributes to output variance but does not directly reveal the input–output functional form, such as local nonlinearity or the detailed structure of interactions. This is why we argue that visual tools such as PDPs and ICE curves provide complementary information to variance decomposition.
Besides GSA, sometimes it is also of interest to understand the behavior of the black-box function when the input of interest is changed. In artificial intelligence and Machine Learning (ML), explainability serves several purposes, including building trust in the model and enhancing transparency, which is especially crucial in fields like medicine [14], and knowledge discovery [15]. Even in engineering, transparency is also vital, as it allows engineers and researchers to verify that the ML model is consistent with established domain knowledge. In this paper, we aim to utilize explainability methods for knowledge discovery in engineering design, viewing the ML model not just as a predictive tool, but also as a source of valuable insights. Recently, there has been growing interest in unifying the frameworks of sensitivity analysis and explainable ML [16].
Techniques from explainable ML and graphical statistics such as Partial Dependence Plot (PDP) [17], Individual Conditional Expectation (ICE) [18], and SHapley Additive exPlanations (SHAP) [19] can be leveraged to understand the behavior of the black-box function. These methods provide insights into how input variables influence the predictive model and interact with each other. Such information is invaluable in engineering and science, helping engineers and scientists extract meaningful insights from the constructed model. The applications of explainable machine learning in engineering are now gaining significant attention, with examples not only in aerospace and mechanical engineering [20, 21, 22, 23], but also in material science [24] and civil engineering [25].
SHAP is arguably one of the most widely used explainability methods. However, despite its widespread adoption, SHAP has certain limitations. While SHAP offers robust insights, the specialized nature of its theoretical foundation may present a learning curve for practitioners not familiar with game theory concepts [26]. Although SHAP decompositions provide faithful local attributions, aggregating those local explanations into a single global summary can conceal heterogeneity across instances [27]. In this context, PDP and ICE offer more intuitive explanations of an ML model due to their simplicity and straightforward depiction of relationships compared to SHAP. They function by altering feature values and observing the corresponding changes in predictions, a concept closely aligned with causal reasoning and easier to understand. Therefore, SHAP is best used together with PDP/ICE or conditional analyses when a clear global picture of feature interactions is required [28].
This paper examines how PDP and ICE curves support engineering design exploration with surrogate models, especially for visualization and global sensitivity analysis. PDPs provide a convenient global sensitivity measure for input importance [29], but their averaging can mask interactions and lead to misleading conclusions if used alone. To address this, we propose to leverage ICE to inspect the individual ICE curves behind the PDP. Even if ICE is still restricted to independent features, it reveals heterogeneous effects and interactions hidden by averaging. We propose a new global sensitivity metric by computing and averaging the standard deviation of each ICE curve, thereby capturing the influence of variable interactions. Furthermore, the standard deviation of importance across all ICE curves is used to quantify the overall impact of interactions, while the ICE correlation value evaluates how such interactions modify the input–output relationship and its nonlinearity. The proposed approach is validated through three case studies: a 5-variable analytical function, a 5-variable wind turbine fatigue problem, and a 9-variable airfoil aerodynamics case.
2 Motivations
The widespread use of high-fidelity simulations in aerospace means that engineers must do more than predict the QoI: they must quantify uncertainty sources and interpret black-box surrogates. This section briefly surveys the convergence of GSA and XAI, and exposes that engineering requires sensitivity analysis that goes beyond classical variance decomposition.
2.1 Global Sensitivity Analysis in Engineering Design
In high-dimensional engineering problems, GSA is essential for knowledge discovery and variable screening; the dominant framework is variance-based sensitivity methods, and are widely considered the “gold standard” in various fields [13, 30]. For high-dimensional problems, estimating variance contributions can be prohibitive in terms of computational time, especially for dummy Monte Carlo implementation [31]. This has led to the popularity of Derivative-based Global Sensitivity Measures (DGSM) [32], which use gradients to bound sensitivity indices, and Leave-one-covariate-out metrics [33], which focus on predictive power rather than variance. Still, many different techniques may alleviate the problem of sensitivity indices computation in high dimensions [34, 35], but with a small number of data points to learn on in high dimensions, these indices remain difficult to estimate [36].
Concerning the computational cost, it is worth noting that many authors have proposed leveraging surrogate models, such as PCE, to perform GSA for complex systems in a tractable and time-efficient manner [30]. Still, when GSA is performed on data-driven surrogates, sensitivity estimates generally mix up input variability and surrogate approximation error unless surrogate uncertainty is explicitly modeled and propagated [30]. Even with many advances, sensitivity indices only provide a single scalar measure of influence, indicating that a variable is important but failing to describe the mechanism of influence. For instance, an index will not reveal how the output behaves across the input’s range (e.g., if the response is monotonically increasing, quadratic, or exhibits a threshold). Understanding this detailed shape of the output response is a prerequisite for making informed design trade-offs [37].
2.2 The Rise of Explainable AI for Engineering
Explainable Artificial Intelligence (XAI) can complement scalar GSA by exposing the functional and instance-level behaviour of complex surrogates used in aerospace design. In practice, XAI delivers two complementary views: local explanations (why a specific prediction was made for a specific input) and global explanations (how the model behaves across the input space). In engineering workflows, local explanations support root-cause analysis and failure investigation, while global views guide design exploration and trade-offs. The most used technique in such a context is SHAP [19] that provides principled local attributions based on cooperative game theory. Importantly, SHAP (local) is distinct from Shapley effects (a global GSA measure): aggregating per-instance SHAP values (e.g., mean absolute SHAP) yields a heuristic global ranking but can mask heterogeneity across the operational design space.
PDP [17] estimates a variable’s marginal profile by averaging the model output over other inputs. While PDP can reveal the shape of a feature’s effect (e.g., linear, nonlinear), it does not directly equate to first-order variance decomposition (such as Sobol’ main effects), because it averages across both main and interaction contributions rather than isolating just the variance attributed to that feature [38]. PDPs are intuitive and widely used to visualise input-output trends, but their averaging hides heterogeneous interactions: when an input has opposite effects in different regions, PDPs can cancel these effects and produce a misleadingly flat profile, a serious hazard in safety-critical cases.
ICE curves [18] plot per-instance response trajectories, revealing heterogeneity that PDP averages obscure. ICE therefore offers richer diagnostic power for aerospace problems, in which understanding is critical. However, their utility in automated engineering design pipelines is hindered by the absence of standardised quantitative metrics or scalar summaries necessary for efficient variable ranking. A key research priority for broader engineering adoption is the development of robust ICE-based scalar metrics, which is the topic of the present paper.
These XAI tools, when used with awareness of their assumptions and computational costs, form a complementary toolkit for interpretable aerospace design.
2.3 Positioning of the Present Work
To address the limitations of the “cancellation effect” in PDPs, Goldstein et al. [18] introduced ICE plots. Instead of plotting the average, ICE plots display the functional relationship for every instance in the dataset, resulting in a set of curves that reveal the full dispersion of model behavior. There is currently a lack of formalised metrics that translate the visual richness of ICE curves into scalar sensitivity indices suitable for automated ranking in engineering design. To that end, this work bridges the gap between the rigorous quantification of variance-based GSA and the visual interpretability of ICE. We propose a novel ICE-based sensitivity metric that quantifies the dispersion of ICE curves. By calculating the standard deviation of the conditional expectations, we derive a metric that:
  1. 1.
    Captures interaction effects that are mathematically canceled out in standard PDP-based metrics.
  2. 2.
    Provides a theoretical lower bound to ICE-based sensitivity.
  3. 3.
    Offers a “correlation” metric to identify whether interactions purely scale the output or fundamentally alter the functional relationship.
This approach allows engineers to retain the intuitive visualization of response curves while ensuring that critical interactions are not masked by averaging artifacts.
3 Background and previous works
3.1 Surrogate model
Let $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{m})^{T}$ be the vector of the input of interest, where $m$ is the number of input variables. Assuming independence between input variables, let us also define the domain of the input variables as $\boldsymbol{\Omega}=\prod_{i=1}^{m}\Omega_{j}$, where $\Omega_{j}$ is the domain for the $i$-th input variable. Let us also define the joint density function $\rho(\boldsymbol{x})=\prod_{j=1}^{m}\rho_{x_{j}}(x_{j})$, where $\rho_{x_{j}}$ where $j=1,\ldots,m$ represents the marginal density. Also, let $y=f(\boldsymbol{x})$ be a black-box function that relates $\boldsymbol{x}$ to the QoI $y$. A surrogate model $\hat{f}(\boldsymbol{x})$ essentially approximates the relationship between $y$ and $\boldsymbol{x}$ using an estimation $\hat{f}(\boldsymbol{x})$, based on the given data set $\mathcal{D}=\{(\boldsymbol{x}^{(i)},y^{(i)})\}_{i=1}^{n}$, where $n$ is the size of the data set. In this paper, we use the data-driven non-intrusive PCE as our surrogate model of choice, although other models could also be applicable.
The PCE model is briefly outlined below, with more detailed explanations available in dedicated literature (e.g.[39]). The expansion reads as
$f(\boldsymbol{x})\approx\hat{f}(\boldsymbol{x})=\sum_{\boldsymbol{\gamma}\in\mathbb{A}_{p}}\beta_{\boldsymbol{\gamma}}\boldsymbol{\Psi}_{\boldsymbol{\gamma}}(\boldsymbol{x}),$ (1)
where the $\boldsymbol{\Psi}_{\boldsymbol{\gamma}}$’s are multidimensional polynomials that are orthonormal with respect to the joint density $\rho(x)$, and $\boldsymbol{\gamma}=\{\gamma_{1},\ldots,\gamma_{m}\}$ with $\gamma_{i}\geq 0$ for each $i$, denotes the individual indices which are members of $\mathbb{A}_{p}$. The coefficients $\boldsymbol{\beta}=\{\beta_{\gamma}\}$ corresponds to the terms of the PCE coefficients, where $P$ is the cardinality of $\mathbb{A}_{p}$.
The index set $\mathbb{A}_{p}$ is generated via the hyperbolic truncation as follows [39]:
$\mathbb{A}_{p}\equiv\{\boldsymbol{\gamma}\in\mathbb{N}^{m}:\|\boldsymbol{\gamma}\|_{q}\leq p\},$ (2)
where
$\|\boldsymbol{\gamma}\|_{q}=\bigg(\sum_{i=1}^{m}\gamma_{i}^{q}\bigg)^{1/q},0<q\leq 1,$ (3)
where $p$ is the polynomial order and $q$ controls the sparsity of the expansion. In this paper, we use $q=1$ since we deal with inputs of medium dimensionality.
A PCE model is constructed by solving $\boldsymbol{F}\boldsymbol{\beta}=\boldsymbol{y}$, where $F_{ij}=\Psi_{j}(\boldsymbol{x}^{(i)})$ forms the $n\times P$ design matrix, with each column representing a polynomial term evaluated at the data points. Here, $\boldsymbol{y}=(y^{(1)},\ldots,y^{(n)})=(f(\boldsymbol{x}^{(1)}),\ldots,f(\boldsymbol{x}^{(n)}))$ represents the responses. In this paper, this system of equations is solved using the least-angle regression algorithm. The optimal order $p$ is sought by exploring various $\mathbb{A}_{p}$ sets, ranging from $\mathbb{A}_{p=1}$ to $\mathbb{A}_{p=p_{max}}$, where $p_{max}$ represents the highest polynomial order considered.
In this study, we employ the data-driven PCE module from UQLab [40], whereby the orthogonal polynomials for the engineering problems are constructed directly from the available data by inferring the probability distributions first.
3.2 Partial Dependence Plots and Individual Conditional Expectations
Let us define $1:m]=\{1,2,\ldots,m\}$. Next, let $S\subset[1:m]$ represent a subset of these indices, and let $C=[1:m]\setminus S$ denote the complementary set of indices not included in $S$. Also, let $\boldsymbol{x}_{S}=(x_{i}|i\in S)$ and $\boldsymbol{x}_{C}=(x_{i}|i\in C)$ represent the vectors of inputs corresponding to $S$ and $C$, respectively. Specifically, we want to build the partial dependence function of the output of the variables indexed by $S$. Let $\Omega_{S}$ be the subdomain for the variables indexed by $S$, i.e., $\boldsymbol{\Omega}_{S}=\prod_{i\in S}\Omega_{i}$. Similarly, $\boldsymbol{\Omega}_{C}=\prod_{i\in C}\Omega_{i}$. Finally, we define the joint probability density functions for these subsets as $\boldsymbol{\rho}_{\boldsymbol{x}_{S}}=\prod_{i\in S}\rho_{x_{i}}(x_{i})$ and $\boldsymbol{\rho}_{\boldsymbol{x}_{C}}=\prod_{i\in C}\rho_{x_{i}}(x_{i})$.
The general formulation of the partial dependence function for $\boldsymbol{x}_{S}$ is written as:
$\hat{f}_{pdp}(\boldsymbol{x}_{S})=\mathbb{E}_{\boldsymbol{x}_{C}}\left[\hat{f}\left(\boldsymbol{x}_{S},\boldsymbol{x}_{C}\right)\right]=\int_{\boldsymbol{\Omega}_{C}}\hat{f}(\boldsymbol{x}_{S},\boldsymbol{x}_{C})\boldsymbol{\rho}_{\boldsymbol{x}_{C}}(\boldsymbol{x}_{C})d\boldsymbol{x}_{C},$ (4)
which represents the marginal expectation of $\hat{f}(\boldsymbol{x})$ over $\boldsymbol{x}_{C}$. For example, if $S=\{1\}$, then we have an interest in building the partial dependence function for the first variable, that is $\hat{f}_{pdp}(x_{1})$. Similarly, if $S=\{1,2\}$, we have an interest in building the joint partial dependence function for $x_{1}$ and $x_{2}$, that is $\hat{f}_{pdp}(x_{1},x_{2})$. PDP essentially shows how a specific set of input variables ($\boldsymbol{x}_{S}$) affects the output of a model, by averaging out the effects of all other variables ($\boldsymbol{x}_{C}$). The partial dependence function can be conveniently calculated using random sampling as follows:
$\hat{f}_{pdp}(\boldsymbol{x}_{S})\approx\frac{1}{n_{mc}}\sum_{i=1}^{n_{mc}}\hat{f}(\boldsymbol{x}_{S},\boldsymbol{x}_{C}^{(i)}),$ (5)
where $n_{mc}$ is the size of the random sampling set and $\boldsymbol{x}_{C}^{(i)}$ is the $i$-th realization of samples drawn from $\boldsymbol{\rho}_{\boldsymbol{x}_{C}}(\boldsymbol{x}_{C})$.
Because PDP averages over other features, it can hide interaction effects; For example, showing a flat line when equal portions of data have positive and negative correlations with the output. This issue is resolved by displaying the ICE curves together with the main PDP [18]. A single ICE curve essentially represents a single $\hat{f}(\boldsymbol{x}_{S},\boldsymbol{x}_{C}^{(i)})$ (see Eq. (5)). That is, a single ICE plot illustrates the effect of changing the input of interest while keeping the other variables constant. To model interaction is to capture non-additivity and more complex effects. Visually, ICE diagnostics reveal the structure of interactions: heterogeneity in curve slopes indicates that the effect of $x_{j}$ depends on $\boldsymbol{x_{C}}$.
Anchoring the PDP and ICE curves at a specific point can better visualize the interactions. To that end, in depicting PDP and ICE curves, we anchor the curves at the centre or mean of the input space. The anchored partial dependence function (i.e., $\hat{f}_{pdp}^{\text{anc}}(\boldsymbol{x}_{S},\boldsymbol{x_{C}})$) is reads as
$\hat{f}_{pdp}^{\text{anc}}(\boldsymbol{x}_{S},\boldsymbol{x_{C}})=\hat{f}_{pdp}(\boldsymbol{x}_{S})-\boldsymbol{1}\hat{f}_{pdp}(\boldsymbol{x}_{S}^{a},\boldsymbol{x}_{C}),$ (6)
where $\boldsymbol{x}_{S}^{a}=\mathbb{E}(\boldsymbol{x}_{S})$ is the anchor point. The ICE curves are also anchored using the same principle. In the discussions and results that follow, we always depict the PDP and ICE curves anchored through this procedure.
4 Derivation of GSA indices from PDP and ICE
A key contribution of this work is to bridge the gap between the visual insight of ICE curves and formal global-sensitivity measures grounded in distributional distances. Below, we make that connection precise, discuss the assumptions, and highlight how our ICE-based metric can serve as a practical proxy for more general sensitivity measures.
Global sensitivity analysis
PDP can provide a global sensitivity derived from the standard deviation of its partial dependence function, i.e., $I_{pdp}$. To make the notation simpler, let us focus on quantifying the importance of a single variable $x_{j}$. Let us first define the variance of the partial dependence function of $x_{j}$ as
$\text{Var}_{x_{j}}\big[\hat{f}_{pdp}(x_{j})\big]=\mathbb{E}_{x_{j}}\left[\left(\hat{f}_{pdp}(x_{j})-\mu_{\hat{f}_{pdp}(x_{j})}\right)^{2}\right],$ (7)
where
$\mu_{\hat{f}_{pdp}(x_{j})}=\mathbb{E}_{x_{j}}\left[\hat{f}_{pdp}(x_{j})\right].$ (8)
The PDP feature importance is then written as [29]:
$I_{pdp}\left(x_{j}\right)=\sqrt{\text{ Var}_{x_{j}}[\hat{f}_{pdp}(x_{j})]}\approx\sqrt{\frac{1}{K-1}\sum_{k=1}^{K}\left(\hat{f}_{pdp}\left(x_{j}^{(k)}\right)-\frac{1}{K}\sum_{k=1}^{K}\hat{f}_{pdp}\left(x_{j}^{(k)}\right)\right)^{2}},$ (9)
where $K$ is the number of samples drawn from $\rho_{x_{j}}(x_{j})$.
$I_{pdp}$ corresponds exactly to the (unnormalised) first‑order effect in the Hoeffding–Sobol (functional ANOVA) decomposition [13]. However, it should not be interpreted as a proportion of total variance unless normalised by the total variance itself, because it only measures the spread of the conditional mean. Moreover, since PDP is computed by averaging model predictions over the complementary variables, its curve can blur interaction effects. This is precisely why ICE curves (and their dispersion) are needed when heterogeneity or interaction is suspected [18, 38].
A similar metric for assessing the interaction term can also be computed. Formally, let $\boldsymbol{x}_{S}=\{x_{i},x_{j}\}$ be the set of two variables that we have an interest in. First, build the joint partial dependence function $\hat{f}_{pdp}(x_{i},x_{j})$. From $\hat{f}_{pdp}(x_{i},x_{j})$, we can calculate $I_{pdp}(x_{i}|x_{j})$ for each value of $x_{j}$, in other words, we compute the partial dependence function of $x_{i}$ on the joint dependence function $\hat{f}_{pdp}(x_{i},x_{j})$. The standard deviation of $I_{pdp}(x_{i}|x_{j})$ is then computed, and the same is done for $I_{pdp}(x_{j},x_{i})$. The standard deviation values for both $I_{pdp}(x_{i}|x_{j})$ and $I_{pdp}(x_{j}|x_{i})$ are then averaged to yield the interaction effect. Regardless, Eq. (9) still represents only the average effect and might cause misunderstanding, which is why we introduce ICE-based feature importance, as explained in the next section.
4.1 ICE-based feature importance
In this paper, we address the limitations of PDP-based feature importance by introducing an ICE-based feature importance metric. While conceptually similar to PDP-based importance, the key distinction is that we first quantify the feature importance for each individual ICE curve before aggregating. Formally, the variance of a single ICE curve with respect to $x_{j}$, given the $i$-th realization of the complementary features in $\Omega_{C}$ (denoted as $\boldsymbol{x}_{C}^{(i)}$), is calculated as follows:
$\mathrm{Var}_{x_{j}}\left[\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)})\right]=\mathbb{E}_{x_{j}}\left[\left(\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)})-\mu_{\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)})}\right)^{2}\right],$ (10)
where the mean of the curve is given by
$\mu_{\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)})}=\mathbb{E}_{x_{j}}\left[\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)})\right].$ (11)
Consequently, the ICE-based importance for a specific instance is defined as the standard deviation of the curve:
$I_{\mathrm{ice}}\left(x_{j};\boldsymbol{x}_{C}^{(i)}\right)=\sqrt{\mathrm{Var}_{x_{j}}\left[\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)})\right]}.$ (12)
In practice, this is approximated empirically using $K$ grid points for feature $x_{j}$:
$I_{\mathrm{ice}}\left(x_{j};\boldsymbol{x}_{C}^{(i)}\right)\approx\sqrt{\frac{1}{K-1}\sum_{k=1}^{K}\left(\hat{f}\left(x_{j}^{(k)},\boldsymbol{x}_{C}^{(i)}\right)-\frac{1}{K}\sum_{l=1}^{K}\hat{f}\left(x_{j}^{(l)},\boldsymbol{x}_{C}^{(i)}\right)\right)^{2}}.$ (13)
This yields a distribution of $I_{\mathrm{ice}}$ values corresponding to different realizations of $\boldsymbol{x}_{C}^{(i)}$. From this distribution, we derive two aggregate metrics:
$\mu_{I_{\mathrm{ice}},x_{j}}=\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C})\right]=\int_{\Omega_{C}}I_{\mathrm{ice}}\left(x_{j};\boldsymbol{x}_{c}\right)\rho_{\boldsymbol{x}_{C}}(\boldsymbol{x}_{c})d\boldsymbol{x}_{c},$ (14)
and
$\sigma_{I_{\mathrm{ice}},x_{j}}^{2}=\mathrm{Var}_{\boldsymbol{x}_{C}}\left[I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C})\right]=\int_{\Omega_{C}}\left(I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{c})-\mu_{I_{\mathrm{ice}},x_{j}}\right)^{2}\rho_{\boldsymbol{x}_{C}}(\boldsymbol{x}_{c})d\boldsymbol{x}_{c}.$ (15)
These two metrics serve distinct but complementary purposes. The first metric, $\mu_{I_{\mathrm{ice}},x_{j}}$, measures global importance. Unlike PDP-based importance, which averages the predictions (often leading to cancellation effects), $\mu_{I_{\mathrm{ice}},x_{j}}$ averages the magnitude of variations, thereby mitigating the issue of heterogeneous effects canceling each other out. Conversely, the second metric, $\sigma_{I_{\mathrm{ice}},x_{j}}$, assesses the influence of interactions. A high dispersion in $I_{\mathrm{ice}}$ indicates that the impact of $x_{j}$ varies significantly depending on the values of the complementary features $\boldsymbol{x}_{C}$, which is a hallmark of strong feature interactions. These metrics can be visualized simultaneously; for example, using a bar plot for $\mu_{I_{\mathrm{ice}},x_{j}}$ with $\sigma_{I_{\mathrm{ice}},x_{j}}$ represented as error bars.
Equation (10) defines the conditional variance of the ICE curve for instance $\boldsymbol{x}_{C}^{(i)}$, and Eq. (12) is its empirical square root. When we then average and compute the spread of $I_{\mathrm{ice}}(x_{j};\boldsymbol{x}_{C}^{(i)})$ across instances, we obtain interpretable quantities in the same units as the model output. One could instead use the conditional variance itself, without taking a square root, as the instance-wise importance measure. That choice has a formal appeal: by the law of total variance, $\mathrm{Var}(Y)=\mathrm{Var}_{X_{j}}\bigl(\mathbb{E}[Y\mid X_{j}]\bigr)+\mathbb{E}_{X_{j}}\bigl[\mathrm{Var}(Y\mid X_{j})\bigr],$ and in classical Sobol’ analysis (assuming input independence), one identifies the first-order component with $\mathrm{Var}_{X_{j}}(\mathbb{E}[Y\mid X_{j}])$ and the complementary term $\mathbb{E}[\mathrm{Var}(Y\mid X_{j})]$ as the residual / interaction-plus component. This decomposition underlies how Sobol’ indices are defined [41]. By contrast, standard deviations are more interpretable (same units as $Y$), but do not decompose additively in the same way. Thus, the two metrics are complementary, and we recommend using the variance for variance decomposition and additivity, while the standard deviation is of better use for interpretability in the original unit of model output.
Recent work has established a rigorous connection between conditional variability summaries and optimal transport-based GSA [42, 43]. Specifically, in [43], the authors propose a sensitivity index based on the squared 2-Wasserstein distance ($W_{2}^{2}$) to quantify the discrepancy between the unconditional model output distribution $\mathcal{L}(Y)$ and the conditional distribution $\mathcal{L}(Y|x_{j})$ [43].
Under the assumption that the distributions are Gaussian or elliptical, this distance admits a closed-form decomposition known as the Wasserstein-Bures semimetric:
$W_{2}^{2}\big(\mathcal{L}(Y),\mathcal{L}(Y|x_{j})\big)=\underbrace{\|\mu_{Y}-\mu_{Y|x_{j}}\|_{2}^{2}}_{\text{Mean Shift}}+\underbrace{\text{Tr}\left(\Sigma_{Y}+\Sigma_{Y|x_{j}}-2(\Sigma_{Y}^{1/2}\Sigma_{Y|x_{j}}\Sigma_{Y}^{1/2})^{1/2}\right)}_{\text{Covariance Shift}}$ (16)
This decomposition highlights two distinct sources of feature importance:
  1. 1.
    The shift in the expected value (first term), which corresponds to the traditional variance-based sensitivity measures [44, 45].
  2. 2.
    The change in the distribution’s shape or dispersion (second term), which captures effects on the output variance and covariance structure.
Our proposed ICE-based metrics align with this theoretical framework by explicitly separating the global average effect (related to the mean shift) from the dispersion of individual ICE curves (related to the covariance shift). This ensures that features influencing the model’s variability, even without shifting the global mean, are correctly identified as important. The key observation is that the partial dependence function $\hat{f}_{pdp}(x_{j})$ averages out the variability of $\hat{f}(x_{j},\boldsymbol{x}_{C})$ across all values of $\boldsymbol{x}_{C}$, potentially smoothing out variations that would otherwise be captured in the ICE curves. Thus, the variability captured by $I_{ice}\left(x_{j};\boldsymbol{x}_{C}^{(i)}\right)$ will always account for more localized variations specific to $\boldsymbol{x}_{C}^{(i)}$, which are not captured by the globally averaged $\hat{f}_{pdp}(x_{j})$. As a result, we demonstrate, in a certain setting, that:
$\mathbb{E}_{x_{C}}\left[I_{ice}(x_{j};x_{C})\right]\geq I_{pdp}(x_{j}).$ (17)
More precisely, we demonstrate that the inequality holds for several classes of functions, most notably for those expressible as a finite expansion of a linear combination of orthogonal bases and functions with some combination of additive and multiplicative separability, as detailed in Appendix A. Since the PCE surrogate model itself constitutes a finite expansion of such bases, the inequality is satisfied when PCE is employed.
4.2 ICE-based correlation values
Beyond quantifying GSA, it is also essential to assess how interactions modify the relationship between inputs and predictions. This enables the identification of variables that change from positive to negative associations (or retain their direction but vary in magnitude) when interacting with others. Such insights are invaluable in design exploration, guiding more informed decisions on controlling design parameters during optimization. To complement the ICE-based feature importance, we introduce the ICE-based correlation values which quantify the correlation between ICE curves and the partial dependence function, written as
$\rho(x_{j},{\boldsymbol{x}_{C}^{(i)}})=\text{Corr}(\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)}),\hat{f}_{pdp}(x_{j})),$ (18)
where Corr denotes the correlation (we use the Pearson correlation coefficient). That is, we compute the correlation coefficient between the ICE curve $\hat{f}(x_{j},\boldsymbol{x}_{C}^{(i)})$ for a specific instance $\boldsymbol{x}_{C}^{(i)}$ and the corresponding partial dependence function $\hat{f}_{pdp}(x_{j}))$ for the variable $x_{j}$.
From this, we define the following quantity:
$\sigma_{\rho_{x_{j}}}=\sqrt{\mathbb{E}_{\boldsymbol{x}_{C}}\left[\left(\rho(x_{j},\boldsymbol{x}_{C})-\mathbb{E}_{\boldsymbol{x_{C}}}\left[\rho(x_{j},\boldsymbol{x}_{C})\right]\right)^{2}\right]},$ (19)
which is essentially the standard deviation of the ICE correlation values, and can be computed from a Monte Carlo simulation. If $\sigma_{\rho}=0$, this indicates that the partial dependence function is similar to ICE curves (in other words, there is no interaction). Conversely, a high value of $\sigma_{\rho}$ suggests that most of the ICE curves differ significantly from the partial dependence line, indicating a strong interaction effect on the trend. Additionally, showing the boxplot of ICE-based correlation values also helps in the assessment.
4.3 First demonstration
The aim of this first demonstration is to investigate the performance of the proposed metrics on a simple problem. At this stage, no surrogate model is used; instead, the methods are applied directly to the actual function. We conducted a comparison with first-order, second-order, and total order Sobol' indices, and averaged SHAP (referred to as $S_{F}$, $S_{T}$, and $\bar{Sh}$, respectively). For the analysis of two-way interactions, we compared the corresponding metrics for two-way interactions, denoted as $I_{pdp,ij}$, $S_{ij}$, and $\bar{Sh}_{ij}$, where $i$ and $j$ represent the indices of the variables, with $i\neq j$. Normalized Shapley effects lie between first-order and total Sobol’ indices [41]. Accordingly, Sobol’ and Shapley derived indices together provide transparent, extensible variance-based diagnostics of input effects and interactions for practical design in engineering.
The three-variable analytical problem is adopted from Goldstein et. al. [18], written as:
$y=0.2x_{1}-5x_{2}+10x_{2}\mathds{1}_{x_{3}\geq 0},$ (20)
where $x_{i}\in[-1,1]$ for $i=1,2,3$. The main challenge with this problem is that $x_{2}$ strongly interacts with $x_{3}$ in such a way that the PDP yields a flat curve (therefore, the $I_{pdp}$ is zero or close to zero). We did not perform a thorough analysis of this problem; instead, we use this problem to check the adequacy of the new metrics (i.e., $\mu_{I,ice}$, $\sigma_{I,ice}$, and $\sigma_{\rho}$). The Sobol' indices and averaged SHAP were calculated using $10^{6}$ samples. Due to the low dimensionality of the problem, the SHAP values are calculated analytically. To calculate the PDP and ICE-based metrics, 10,000 ICE curves were drawn from the actual function but the SMT-Explainability toolbox allows us to do the same based on a surrogate model [46, 47].
(a) $x_{1}$
(b) $x_{2}$
(c) $x_{3}$
Figure 1: PDP and ICE plots for the three-variable problem.
The centered PDP and ICE plots for this problem are shown in Fig. 1. It is evident that the influence of $x_{2}$ on $y$ varies depending on the value of $x_{3}$, and vice versa. As a result, the PDPs for $x_{2}$ and $x_{3}$ appear as almost flat lines, leading to a near-zero PDP-based feature importance for both variables. However, this is misleading, as the true importance of these variables is not zero. On the other hand, $x_{1}$ does not interact with $x_{2}$ and $x_{3}$, which is why its PDP and ICE curves are essentially identical. Moreover, its overall impact is actually less significant than that of $x_{2}$ and $x_{3}$.
The complete GSA results are presented in Table 1. The first-order Sobol' indices are essentially zero for all variables. However, the total order Sobol' indices are close to 1 for $x_{2}$ and $x_{3}$, suggesting that the only contribution to the total variance comes from the interaction between $x_{2}$ and $x_{3}$. As anticipated, $I_{pdp}$ does not suggest that $x_{2}$ and $x_{3}$ are more significant than $x_{1}$, due to the previously mentioned challenge. In contrast, Averaged SHAP accurately identifies the importance of $x_{2}$ and $x_{3}$. Unlike $I_{pdp}$, $\bar{Sh}$ is based on the impact of each feature on the model’s predictions and is capable of detecting subtle yet crucial relationships that $I_{pdp}$ might overlook. It can also be observed that ICE-based feature importance yields finite and reasonable values for $x_{2}$ and $x_{3}$. Notably, $\mu_{I,ice}$ effectively identifies the importance of $x_{2}$ and $x_{3}$ by focusing on the importance of each ICE curve. Thus, it can be concluded that $\mu_{I,ice}$ addresses the main limitation of $I_{pdp}$. The $\sigma_{I,ice}$ value that equals 0 for $x_{1}$ is the consequence of its non-interacting nature. However, $\sigma_{I,ice}=0$ could also mean another thing. In fact, a zero value of $\sigma_{I_{\mathrm{ice}}}$ means the per-instance ICE spread is identical for all conditioning values; this can indicate either (i) a genuinely non-interacting additive input, as in our example, or (ii) the presence of interactions that alter curve shape (e.g. flip sign or slope) while preserving the integrated spread. Hence, in this second case, visual inspection of ICE curves or slope-based diagnostics is required to disambiguate these cases, especially when relying on surrogate models. For the case of $x_{2}$, this means that the importance magnitude of each ICE curve of $x_{2}$ is the same, although the gradient shifts depending on how it interacts with $x_{3}$. Lastly, the finite value of $\sigma_{I,ice}$ for $x_{3}$ is reasonable because, as can be observed in Fig. 1(c), each ICE curve yields different $I_{ice}$.
Lastly, $\sigma_{\rho}$ is not a measure of GSA but a change in the trend of ICE curves with respect to the PDP. The fact that $\sigma_{\rho}=0$ for $x_{1}$ suggests that $x_{1}$ does not interact with other variables. Furthermore, the values of $\sigma_{\rho}=1$ for $x_{2}$ and $x_{3}$ indicates strong impact of interactions on the trend. As suggested by Fig. 1 again, the correlations between ICE curves and PDP for $x_{2}$ are either -1 or 1. Although the finite correlations result from the calculation process and should theoretically be zero when correlating with a flat function, this nonetheless highlights a subtle shift in the trend. As for $x_{3}$, the sharp discontinuity is the reason why the standard deviation of the ICE correlation values is non-zero.
Table 1: GSA Results for the 3-variable problem.
Variable $x_{1}$ $x_{2}$ $x_{3}$
$S_{F}$ 0.001 0.004 0.001
$S_{T}$ 0.003 0.997 0.998
$\bar{Sh}$ 0.552 1.750 0.750
$I_{pdp}$ 0.116 0.003 0.010
$\mu_{I,ice}$ 0.116 2.908 2.505
$\sigma_{I,ice}$ 0.000 0.000 1.449
$\sigma_{\rho}$ 0.000 1.000 1.000
The two-way GSA results are presented in Table 2. It is noteworthy that all metrics (namely, two-way $I_{pdp}$, two-way $\bar{Sh}_{ij}$, and second-order Sobol' indices) successfully identify the interaction between $x_{2}$ and $x_{3}$. It is an expected behavior as Second-order Sobol' indices quantify how much the combined interaction of two input variables contributes to the total output variance, over and above the sum of their individual effects. In the case of $\mu_{I,ice}$, this interaction information is incorporated into $\mu_{I,ice}$ and complements it. Specifically, $\mu_{I,ice}$ and its corresponding $\sigma_{I,ice}$ reflect the overall impact of the interaction, while $I_{pdp,ij}$ quantifies the influence of specific interactions.
Table 2: Two-way GSA Results for the 3-variable problem.
Variable $x_{1}-x_{2}$ $x_{1}-x_{3}$ $x_{2}-x_{3}$
$I_{pdp,ij}$ 0.000 0.000 0.771
$S_{ij}$ 0.004 0.004 0.993
$\bar{Sh}_{ij}$ 0.000 0.000 1.246
5 Demonstration on analytical and engineering problems
Metrics for analytical problems were evaluated on the original function, while surrogate models were used for engineering cases where the true function was unknown. For all problems, we conducted a comparison with other sensitivity metrics (e.g., Sobol' indices and averaged SHAP).
5.1 5-variable Friedman function
We first demonstrated and compared the capabilities of ICE and PDP on the 5-variable Friedman function [48], written as:
$f(\boldsymbol{x})=10\text{ sin}(\pi x_{1}x_{2})+20(x_{3}-0.5)^{2}+10x_{4}-5x_{5}$ (21)
where $x_{i}\in[0,1]$, for $i=1,\ldots,5$. The equation shows interaction only between $x_{1}$ and $x_{2}$; thus, explainability methods should be able to correctly capture this relationship.
5.1.1 Global sensitivity analysis
First, let us compute the GSA values from Sobol' indices, averaged SHAP, PDP, and ICE. The comparison of various GSA metrics is presented in Table 3. Notably, all metrics produce the same ranking of importance, with the most significant variable being $x_{4}$ followed by $x_{1}$ and $x_{2}$, with $x_{3}$ and $x_{5}$ are considered to be less significant. Interestingly, $I_{pdp}$ and $\mu_{I,ice}$ provide identical importance values for $x_{3}$, $x_{4}$, and $x_{5}$. This occurs because the centered ICE curves for variables without interactions are essentially the PDP itself, leading to the mean of $I_{ice}$ (i.e., $I_{pdp}$) being identical to $I_{pdp}$. Conversely, the $\mu_{I,ice}$ values for $x_{1}$ and $x_{2}$ are higher than their $I_{pdp}$ counterparts, which is due to the interaction effects considered in the calculation of $\mu_{I,ice}$. In this regard, if the value $I_{pdp}$ equals $\mu_{I,ice}$, this signals that the interaction is not present. Conversely, when $\mu_{I,ice}>I_{pdp}$, this signals that the interaction term is present since this means that the ICE curves are not exactly similar to the PDP in terms of the trend and impact. Note that this approach is similar to interpreting first-order and total order Sobol' indices. It is also evident in this problem that the values of $S_{T}$ and $S_{F}$ are identical for $x_{3}$, $x_{4}$, and $x_{5}$, while $S_{T}$ for $x_{1}$ and $x_{2}$ are higher than $S_{F}$ for $x_{1}$ and $x_{2}$.
Table 3: GSA results for the Friedman problem.
Variable $x_{1}$ $x_{2}$ $x_{3}$ $x_{4}$ $x_{5}$
$S_{F}$ 0.195 0.198 0.091 0.348 0.086
$S_{T}$ 0.272 0.271 0.091 0.348 0.086
$\bar{Sh}$ 2.278 2.278 1.707 2.503 1.256
$I_{pdp}$ 2.176 2.175 1.509 2.908 1.454
$\mu_{I_{ice}}$ 2.415 2.422 1.509 2.908 1.454
$\sigma_{I_{ice}}$ 0.856 0.852 0.000 0.000 0.000
$\sigma_{\rho}$ 0.110 0.114 0.000 0.000 0.000
Table 4: Two-way GSA results for the Friedman problem. The other interactions yield essentially zero importance.
Variable $x_{1}-x_{2}$ $x_{1}-x_{4}$ $x_{2}-x_{4}$
$I_{pdp,ij}$ 0.995 0.000 0.000
$\bar{Sh}_{ij}$ 0.585 0.000 0.000
$S_{ij}$ 0.076 0.001 0.001
The barplot of $I_{pdp}$, $\mu_{I,ice}$, and $\bar{Sh}$ for the Friedman regression problem is further shown in Fig. 2(a). Here, it can be seen that $\mu_{I,ice}$, with the errorbar showing $\mu_{I,ice}\pm\sigma_{I,ice}$ provides more information than $I_{pdp}$ and $\bar{Sh}$. Again, a large vertical bar indicates a stronger effect of interactions, as evidenced in $x_{1}$ and $x_{2}$, but non-existent in other variables. Further, Fig. 2(b) depicts the boxplots of the correlation between the ICE curves and PDP. In this figure (and similar figures throughout the paper), the lower and the upper bounds of the box show the first ($Q_{1}$) and the third quartile ($Q_{3}$), respectively, the solid line within the box shows the median ($Q_{2}$), the lower and the upper“whisker” show $Q_{1}-1.5(Q_{3}-Q_{1})$ and $Q_{3}+1.5(Q_{3}-Q_{1})$, and the crosses show individual data that lie outside the “whiskers”. For $x_{1}$ and $x_{2}$, the overall correlations are positive, but some ICE curves deviate significantly from the PDP (indicating strong interaction), particularly those with lower correlation values. The non-zero values of $\sigma_{\rho}$ also signify that the effect of the interaction changes the trend of the ICE curves. It is worth noting that a non-zero value of the correlation between the ICE curve and PDP indicates a change in nonlinearity or trend. The $\sigma_{\rho}$ value for $x_{3}$, $x_{4}$, and $x_{5}$ is exactly zero because the trend for the ICE curve is exactly similar to its PDP. Let us now analyze the interaction GSA-values as shown in Table 4. The two-way PDP feature importance, SHAP, and second-order Sobol' indices (see Table 4). further confirm the presence of the interaction between $x_{1}$ and $x_{2}$. In this regard, Second-order Sobol indices quantify the interaction effects between pairs of input variables on the output variance, while interaction SHAP values measure the specific contribution of interactions between inputs to a model’s prediction. It was observed that the estimated interaction metrics are zero or nearly zero for all cross terms but $x_{1}-x_{2}$. Note that this information should be digested together with an analysis of the visualization of PDP and ICE, as shown next.
(a) Barplot of $I_{pdp}$, $\mu_{I,ice}$, and $\bar{Sh}$.
(b) Boxplot of ICE-based correlation values.
Figure 2: Barplot of some GSA metrics and boxplot of ICE correlation values for the Friedman regression problem.
5.1.2 Visualization
The first variable of interest is $x_{1}$, with the centered PDP/ICE and SHAP plots shown in Figs. 3 and 4, respectively. The plots for $x_{2}$ are similar, as both variables appear only in the term $10\sin(\pi x_{1}x_{2})$. The curves are colored by $x_{2}$ and $x_{4}$, where $x_{1}$ interacts solely with $x_{2}$. Notably, the centered PDP/ICE and SHAP plots reveal distinct insights: the effect of $x_{1}$ varies with the value of $x_{2}$, reflecting their interaction. For instance, when $x_{2}$ is small, the impact of $x_{1}$ is minimal, resulting in an almost flat ICE curve. However, as $x_{2}$ increases, the impact of $x_{1}$ becomes more significant, with the overall trend first increasing and then decreasing (which is due to the sinusoidal term). This effect is quantified and shown by the non-zero values of $\sigma_{I,ice}$ and $\sigma_{\rho}$ discussed earlier. Further, it can be seen that the coloring seems to be disordered when the ICE curves for $x_{1}$ are color-coded according to $x_{4}$; which indicates that there is no interaction between how the change in $x_{4}$ affects $x_{1}$.
(a) PDP and ICE
(b) PDP and ICE
Figure 3: PDP and ICE plots of $x_{1}$ for the 5-dimensional Friedman regression problem.
The SHAP plots for $x_{1}$, while they do reveal interactions similar to the PDP/ICE plots, do not make it immediately clear that the effects of $x_{1}$ and $x_{2}$ are relatively flat for small values of both variables. This is because SHAP values emphasize the contribution of input variables to individual predictions, rather than illustrating the global effect of changing an input variable. As a result, while SHAP is valuable on its own, certain aspects may not be easily interpreted from it.
(a) SHAP
(b) SHAP
Figure 4: SHAP dependence plots of $x_{1}$ for the 5-dimensional Friedman regression problem.
The joint PDP and SHAP interaction plots for $x_{1}-x_{2}$ are shown in Fig. 5, respectively (the other interactions are not shown since they are essentially non-existent). Note that the SHAP visualization was generated by holding the values of the other variables constant at the center of the input space. The color of those two plots corresponds to the value of partial dependence, for the left PDP interaction plot, and to the values of SHAP for the right interaction SHAP plot. It is worth noting that both plots visualize two different aspects. That is, the joint PDP plot shows the average impact of changing $x_{1}$ and $x_{2}$ together on the Friedman function. Such a plot is primarily useful if the aim is to see both the main effects and the combined impact. Conversely, the SHAP interaction plot isolates the specific interaction between these two variables, allowing us to focus solely on the synergistic effect that cannot be attributed to either variable independently.
(a) Joint partial dependence function
(b) Interaction SHAP
Figure 5: Two-way dependence plots of $x_{1}-x_{2}$ for the Friedman function according to the joint partial dependence function and interaction SHAP.
5.2 5-variable wind turbine fatigue problem
We now address the uncertainty analysis of a wind turbine fatigue problem, where the uncertainties involve five independent random variables related to wind conditions [49]. The input variables considered are wind speed ($V_{hub}$), wind direction ($\theta_{w}$), wave height ($H_{s}$), wave period ($T_{p}$), and wind-platform direction ($\theta_{p}$) with their histograms are shown in Fig. 6. Notice that the data was normalized within $1,1]^{5}$, as provided by the original paper. The side-side tower base bending moment ($M_{x,twr}$) is selected as the output of interest due to its nonlinear behavior, evaluated using FAST, NREL’s aeroelastic code for simulating the coupled dynamic response of wind turbines [50]. Subsequently, a PCE model with $p_{max}=5$, $q=1$, and data-driven orthogonal polynomials was built using 2000 samples, yielding $R^{2}=0.97$ which is more than sufficient for knowledge discovery purposes.
Figure 6: Histograms of the input variables for the 5-variable wind turbine fatigue problem.
5.2.1 Global Sensitivity Analysis
Fig. 7(a) shows the GSA results of three metrics ($I_{pdp}$, $\mu_{I,ice}$, and $\bar{Sh}$) for the 5-variable wind turbine fatigue problem (the Sobol' indices are not shown due to the different scale, but instead shown in Fig. 8). All GSA metrics agree that the most important variable is $V_{hub}$ followed by $\theta_{w}$ and $H_{s}$. On the other hand, $T_{p}$ and $\theta_{p}$ are deemed to be insignificant. The averaged SHAP perceives the importance of $\theta_{w}$ is about half that of $H_{s}$, with $I_{pdp}$ and $\mu_{I,Ice}$ perceiving that the former is slightly important than the latter. It is interesting to see from the barplot that, although $V_{hub}$ is the most impactful variable, the effect of interaction is felt the strongest by $\theta_{w}$ and $H_{s}$ as indicated by the higher value of $\sigma_{I,ice}$ compared to $V_{hub}$. The cause for this will be discussed soon by visualizing the PDP and ICE curves. It is worth noting again that $\sigma_{I,ice}$ does not indicate which combinations of variables contribute the strongest to the interaction; however, the information from $\sigma_{I,ice}$ is useful in revealing the variability of the impact, complemented by the average impact from $\mu_{I,ice}$. Lastly, the impact of interactions on $T_{p}$ and $\theta_{p}$ is nevertheless small, and their importance can be neglected.
The boxplots of ICE-based correlation values are shown in Fig. 7(b). The actual $\sigma_{\rho}$ values for $V_{hub}$, $\theta_{w}$, $H_{s}$, $T_{p}$, and $\theta_{p}$ are $0.058$, $0.275$, $0.160$, $0.444$, and $0.573$, respectively. These boxplots suggest that the interaction does not significantly alter the main trend of how $V_{hub}$ influences the moment. However, a notable shift in trends is evident for $\theta_{w}$ and $H_{s}$, indicating that the marginal behavior of these two variables is strongly influenced by their interactions with each other or with other variables. While the shifts in trend due to $T_{p}$ and $\theta_{p}$ appear significant, it is important to note that these changes may be the result of minor variations, as the impact of these two variables is relatively insignificant. Finally, quantification of the impact of interactions is shown in Fig. 8, with the diagonal entries showing the main effect (for Sobol' indices, the first-order Sobol' indices). All GSA agree that the most important interaction is between $H_{s}$ and $\theta_{w}$, followed by $V_{hub}-H_{s}$. Notice that the off-diagonal colors representing the correlations between $V_{hub}$, $\theta_{w}$, and $H_{s}$ are brighter for averaged SHAP and $I_{pdp}$ compared to Sobol' indices. This difference arises because Sobol' indices are variance-based, which amplifies the importance of the most influential variables.
(a) Barplot of $I_{pdp}$, $\mu_{I,ice}$, and $\bar{Sh}$
(b) Boxplot of ICE correlation values
Figure 7: GSA results for the 5-variable wind turbine fatigue problem ($I_{pdp}$, $\mu_{I,ice}$, and $\bar{Sh}$) and the ICE correlation values.
(a) Sobol' indices
(b) $I_{pdp}$
(c) SHAP
Figure 8: Two-way GSA results for the 5-variable wind turbine fatigue problem. The diagonal entries show the main effect importance, while the off-diagonal entries show the two-way interaction importance.
5.2.2 Visualization
Fig. 9 shows the PDPs and ICE curves for the three most significant variables, clearly showing how they affect the model’s behavior. The effect of $V_{hub}$ is nonlinear, with $M_{x,twr}$ generally increasing as $V_{hub}$ rises, although there is a slight decrease when $V_{hub}$ slightly exceeds its nominal value. Similarly, an increase in $\theta_{w}$ tends to raise $M_{x,twr}$. However, the actual trend influenced by $H_{s}$ is more complex, as revealed by the ICE plots and two-way PDPs. The harmonic looking pattern in $M_{x,twr}$ due to changes in $H_{s}$ is evident, with peaks observed around $H_{s}=-0.7$ and $H_{s}=0.7$. Additionally, as can be seen from the ICE plots, the impact on the bending moment is more pronounced when $\theta_{w}$ is large. Linking this observation with the GSA metrics, the large range of $\hat{f}_{pdp}$ of $V_{hub}$ contributes to its large importance. However, the impact of interaction itself is relatively small, as can be seen from the smaller deviation of ICE curves from the PDP. On the other hand, the deviation of ICE curves from their partial dependence function can be clearly seen for $\theta_{w}$ and $H_{s}$, particularly the latter; Overall, this contributes to their relatively large values of $\sigma_{I,ice}$. For comparison, Fig. 10 shows the SHAP dependence plots for the same set of variables. While the insights from SHAP and PDP/ICE are similar, the former appears more disordered, which is understandable given that SHAP values exist for every point in the input space. In that sense, it is easier to understand the effect of changing $V_{hub}$, $\theta_{w}$, and $H_{s}$ since PDP and ICE curves are smooth.
(a) $V_{hub}$
(b) $\theta_{h}$
(c) $H_{s}$
Figure 9: PDP and ICE plots for the first three important variables.
(a) $V_{hub}$
(b) $\theta_{h}$
(c) $H_{s}$
Figure 10: SHAP dependence plots for the first three important variables.
Fig. 11 depicts the two-way PDP plots for the combinations of the three most important variables (i.e., $V_{hub}-\theta_{w}$, $V_{hub}-H_{s}$, and $\theta_{w}-H_{s}$). The complex interactions between $\theta_{w}$ and $H_{s}$ are clearer when plotting their joint partial dependence function. For example, it can be observed that the marginal impact of $H_{s}$ becomes larger when the corresponding $\theta_{w}$ is large. To ensure the reliability of wind turbines, the GSA metrics and plots indicate that specific conditions can result in significant bending moments. As previously mentioned, these conditions occur when $\theta_{w}$ and $V_{hub}$ are large, and $H_{s}$ is approximately 0.7-0.7 or $0.7$. One limitation of the joint partial dependence plot is that it can be challenging to discern the interaction strength between $V_{hub}$ and $\theta_{w}$ (which is actually small). Therefore, this information should be considered alongside the analysis of GSA metrics for a more complete understanding.
(a) $V_{hub}-\theta_{h}$
(b) $V_{hub}-H_{s}$
(c) $\theta_{h}-H_{s}$
Figure 11: Two-way PDP plots for the first three important variables.
5.3 9-variable Airfoil database problem
5.3.1 Data set
We evaluated the usefulness of PDP and ICE, particularly the global sensitivity metrics derived from ICE, for knowledge discovery in an airfoil database from Agarwal et al. [51], where aerodynamic coefficients were computed using OpenFOAM at $Re=10^{5}$. The dataset was built by fitting eight CST parameters [52] to airfoils from the UIUC Airfoil Data Site. Aerodynamic performance was then obtained using the simpleFOAM solver with the viscous Spalart–Allmaras turbulence model. Simulations were conducted at 13 different angles of attack ($\alpha$), spanning from -4 to 8 degrees, with an increment of 1 degree. The output of interest is the drag coefficient ($C_{d}$). Four basis functions are deemed sufficient for this purpose (Fig. 12(a)). Hence, in total there are nine input variables, collected into a vector $\boldsymbol{x}=[\alpha,A_{l,1},A_{l,2},A_{l,3},A_{l,4},A_{u,1},A_{u,2},A_{u,3},A_{u,4}]^{T}$, where $A$ is the CST coefficients and the subscript $l$ and $u$ indicate upper and lower surface, respectively.
After pre-processing to remove erroneous aerodynamic data, particularly abnormal $C_{d}$$\alpha$ trends, the dataset was filtered, yielding 6,630 samples spanning 510 airfoils at 13 angles of attack. A snapshot of 100 airfoils with the airfoil created from the centre of the CST coefficients’ range (see Table 5) is depicted in Fig. 12(b). A data-driven PCE model with $p_{max}=4$ and $q=1$ was then constructed using 6,000 data points, with the remaining 630 points set aside for testing. The model achieved a coefficient of determination of 0.994, indicating a high level of accuracy suitable for knowledge discovery purposes.
Variable Lower Bound Upper Bound
$\alpha$ [deg] -4.000 8.000
$A_{l,1}$ -0.236 -0.077
$A_{l,2}$ -0.249 0.090
$A_{l,3}$ -0.379 0.068
$A_{l,4}$ -0.219 0.237
$A_{u,1}$ 0.131 0.338
$A_{u,2}$ 0.104 0.475
$A_{u,3}$ 0.081 0.417
$A_{u,4}$ 0.041 0.429
Table 5: Variables and their corresponding lower and upper bounds for the 9-variable airfoil problem.
(a) Basis functions
(b) CST-generated airfoils
Figure 12: Four basis functions of CST and a snapshot of airfoils from the database with the airfoil at the centre of the CST range.
5.3.2 Result
The GSA and ICE correlation results for the 9-variable airfoil problem are shown in Fig. 13. The two-way GSA results are shown in Fig. 14 (refer to the off-diagonal entries). Fig. 15 shows the one-dimensional PDP and ICE plots for select variables. As can be seen in Fig. 13, all GSA metrics agreed on ranking the importance of the four first important input variables (i.e., $\alpha$, $A_{u,2}$, $A_{u,3}$, and $A_{u,1}$). From GSA, it can be seen that $\alpha$ is the most important parameter in terms of drag generation, which makes sense since changing $\alpha$ significantly changes the pressure and shear-stress distribution at the airfoil; However, our main focus here is on the CST parameters. All GSA metrics consistently indicate that the upper surface CST parameters dominate drag generation, reflecting their influence on boundary-layer separation. The most critical variable is $A_{u,2}$, which shapes the upper surface near the maximum thickness. Minor discrepancies in the ranking of less influential variables (lower-surface terms and $A_{u,4}$) are negligible.
It can also be seen that the values of $\mu_{I,ice}$ are larger than $I_{pdp}$. In this regard, such differences are notably relatively significant for the first three lower surface CST variables, indicating that the respective $I_{pdp}$ values are smaller than $\mu_{I,ice}$ due to alternating positive and negative correlations (hence, an almost flat PDP, discussed next). It can also be seen that the impact of interactions between variables on the ICE curves is strong, as evidenced by the significantly high value of $\sigma_{I,ice}$ for all variables (see the error bar in Fig. 13). Therefore, the drag response surface is characterized by nonlinearity and strong interactions between input variables. It is worth noting that the values of $\sigma_{I,ice}$ themselves do not directly inform which variables contribute to the interactions for a specific input variable. Although it makes sense that the interaction is primarily due to $\alpha$ since it is the most important input variable, the interaction with other CST variables also contributes to such differences.
(a) GSA results
(b) ICE correlation values
Figure 13: GSA and ICE correlation values results for the 9-variable airfoil problem
The two-way GSA results are shown in the off-diagonal entries of Fig. 14. We focus exclusively on the CST coefficients, omitting the impact of the angle of attack to enhance plot clarity. Compared to the main effects, as expected, the interaction effects are relatively weaker. As discussed, the strongest interaction occurs between $A_{u,2}$ and $A_{u,3}$, which is expected since both are the most influential variables. However, it is important to note that these interactions alter the relationships between the input variables and the drag coefficient.
(a) Sobol' indices
(b) $I_{pdp}$
(c) SHAP
Figure 14: Two-way GSA results for the 9-variable airfoil database problem. The diagonal entries show the main effect importance, while the off-diagonal entries show the two-way interaction importance.
The boxplot of ICE correlation values presented in Fig. 13(b) offers several key insights. First, it is evident that lower surface CST parameters, particularly $A_{l,1}$ and $A_{l,4}$, exhibit high dispersion in terms of ICE correlation values, indicating significant variations in their ICE curves due to interactions with other variables (as confirmed in Fig. 15). Conversely, the lower dispersion for upper surface CST parameters suggests that interactions have a minimal effect on the overall trend, though the magnitude of the impact may still vary, as seen in the corresponding PDP and ICE plots. Overall, the lower surface CST parameters are more prone to trend changes due to interactions. In contrast, the upper surface parameters show less variation in trend but remain the most important variables in terms of significance. For the PDP and ICE curves of all variables, the trend consistently remains monotonic, reinforcing the hypothesis that the drag response of an airfoil is an unimodal function [53], even in the presence of nonlinearity.
The impact of $A_{l,1}$ is characterized by complex interaction with $A_{u,1}$ (see Fig. 15(a)). In this sense, a nonlinear trend that also changes in terms of correlation with the value of $A_{l,1}$ is present. Such an alternating trend results in a partial dependence function close to the flat line, which is why its $I_{pdp}$ is notably smaller than $\mu_{I,ice}$. The same trend also exists for $A_{l,4}$ as shown in Fig. 15(b). We decided to plot how $A_{l,4}$ interacts with $\alpha$ for demonstration purposes. In this regard, we can see that the association of the PDP of $A_{l,4}$ with its input value changes depending on the angle of attack. To be more exact, increasing $A_{l,4}$ results in higher drag for a positive angle of attack, while the opposite trend occurs for a smaller angle of attack. The PDP and ICE curves for $A_{u,2}$ clearly demonstrate its interaction with $A_{u,3}$ in influencing drag (see Fig. 15(c)). Notably, the effect of $A_{u,2}$ is most pronounced when $A_{u,3}$ is at a high value. Conversely, this implies that altering $A_{u,2}$ has minimal impact on drag when $A_{u,3}$ is small. Aerodynamically, the combinations of small $A_{u,2}$ and $A_{u,3}$ correspond to airfoils with reduced thickness, resulting in lower drag.
The SHAP dependence plots roughly reveal the same trend as the PDP and ICE plots (see Fig. 16). However, the colouring of SHAP plots might be too crowded, especially if the dots are too scattered and complex interactions exist. For example, it is quite difficult to decipher the trend in the $A_{u,2}$ plot regarding how it interacts with $A_{u,3}$. In this regard, PDP and ICE provide more straightforward and easier-to-read plots than SHAP.
(a) $A_{l,1}$
(b) $A_{l,4}$
(c) $A_{u,2}$
Figure 15: PDP and ICE plots for selected inputs for the 9-variable airfoil database problem
(a) $\alpha$
(b) $\alpha$
(c) $\alpha$
Figure 16: SHAP dependence plots for selected inputs for the 9-variable airfoil database problem
Finally, Fig. 17 visualizes the joint dependence plot for select variables, namely: $A_{l,1}-A_{u,1}$, $A_{l,4}-\alpha$, and $A_{u,2}-A_{u,3}$. The trend of $A_{u,2}-A_{u,3}$, in which it can be seen that minimum drag is achieved by decreasing the two variables together (i.e., due to the reduced thickness). On the other hand, the trend is more complex and wavy for the joint plot for $A_{u,1}$ and $A_{l,1}$. The change in the association of $A_{l,1}$ to drag as a function of $A_{u,1}$ can also be seen from this plot.
(a) $A_{l,1}-A_{u,1}$
(b) $A_{l,4}-\alpha$
(c) $A_{u,2}-A_{u,3}$
Figure 17: PDP and ICE plots for selected inputs for the 9-variable airfoil database problem
5.4 Interim summary
In the workflow of knowledge discovery in engineering, our ICE-based summaries $\mu_{I_{\mathrm{ice}},x_{j}}$ and $\sigma_{I_{\mathrm{ice}},x_{j}}$ are useful complements, since they capture instance-level spread and heterogeneity that scalar variance indices aggregate away. Based on the experiments that we performed, we summarize our main findings as follows:
  1. 1.
    The ICE-based feature importance ($\mu_{I,ice}$ with $\sigma_{I,ice}$) provides richer insights than the traditional PDP-based metric ($I_{pdp}$). While $\mu_{I,ice}$ captures effects missed by $I_{pdp}$, $\sigma_{I,ice}$ quantifies interaction influence. Complementing SHAP and Sobol’ indices, the ICE-based approach offers a distinct and informative perspective.
  2. 2.
    The value of $\sigma_{I,ice}$ reflects only the overall interaction impact, not which variables are involved. Interacting variables can be identified using joint PDPs, two-way Sobol’ indices, or two-way SHAP values. Combining these visualizations with GSA metrics ($\mu_{I,ice}$, $\sigma_{I,ice}$, and two-way GSA) enables a more comprehensive understanding of the model’s behavior.
  3. 3.
    The ICE correlation values provide a quantitative view of how interactions modify input–output relationships, revealing changes in linearity or correlation. However, they should be interpreted with caution, as they indicate association rather than the magnitude of impact.
  4. 4.
    PDP (with ICE) and SHAP visualizations generally convey similar insights, though PDP and ICE are often more intuitive as they depict smooth model slices and averages. Joint dependence and SHAP interaction plots, while distinct, offer complementary perspectives. Combined with the proposed GSA metrics, these visualizations enable a comprehensive understanding of the predictive model.
Our findings highlight the utility of ICE-based feature importance. While $\mu_{I,ice}$ offers a more comprehensive view of the input variable’s impact by capturing effects that $I_{pdp}$ might miss, $\sigma_{I,ice}$ further enriches the analysis by revealing how interactions with other variables influence this impact.
6 Conclusion and future works
The primary goal of this paper is to evaluate the effectiveness of PDP and ICE in supporting engineering design exploration and analysis. More importantly, we introduce new metrics that enhance the insights obtainable from existing PDP- and ICE-based approaches. Specifically, we propose an ICE-based feature importance metric that addresses the alternating-association issue inherent in PDP-based importance. The key idea is to compute the mean and standard deviation of the feature importance derived from each ICE curve, providing both a global average effect and information on how interactions with other variables influence a feature’s significance. This allows users to assess not only the average impact of an input variable but also its potential strongest or weakest influence on the predictive model. The standard deviation summarizes the effect of higher-order interactions in a single value. Additionally, we assess the correlation between ICE curves and PDP to further examine how interactions modify the input–output relationship.
We applied PDP and ICE to a 5-variable Friedman function, a 5-variable wind-turbine fatigue problem, and a 9-variable airfoil dataset to demonstrate the insights enabled by these methods and the new sensitivity metrics. PDP and ICE reveal information not easily obtained from SHAP, and each method contributes a distinct perspective. The proposed metrics further provide quantitative interaction-aware sensitivity measures that strengthen knowledge discovery in engineering design analysis.
Future work may explore integrating PDP and ICE insights into optimization frameworks, using the proposed sensitivity metrics to guide design decisions by prioritizing influential variables. Advanced visualization techniques that dynamically reveal interactions captured by ICE curves and PDPs would further improve interpretability, especially in complex, nonlinear systems. The standard-deviation–based metrics could also support confidence-interval visualization for PDP, though a full uncertainty-quantification pipeline is needed to incorporate additional sources of uncertainty, including surrogate-model approximation errors. Developing such a pipeline would provide a more complete view of uncertainty in high-fidelity engineering simulations. Finally, automating insight extraction from PDP, ICE, and the proposed metrics could reduce reliance on expert judgment and improve usability in real-world engineering workflows.
Conflict of interest
The authors have no Conflict of interest to declare.
Data availability and replication
The data on which this work is based is available on GitHub: https://github.com/kanakaero/Dataset-of-Aerodynamic-and-Geometric-Coefficients-of-Airfoils (the data do not belong to the authors).
Both the produced data and the code developed to reproduce the results on aerodynamics and airfoil problems will be available upon request.
Appendix A Proving Inequalities Between PDP and ICE-based global sensitivity metrics
In this section, we prove an inequality involving the PD Function and the ICE for a large class of functions. Below, let $1\leq j\leq m$ and let $C=\{1,2,\ldots,m\}\setminus\{j\}$. We have an interest in the GSA metrics for a single variable indexed by $j$ (i.e., $x_{j}$), with $\boldsymbol{x}_{C}$ referring to other variables indexed in $C$. Also, let $\mathbb{E}[\cdot]$ and $\mathbb{V}[\cdot]$ denote the expectation and the variance operator, respectively, and the subscript denotes that the operation is being taken with respect to a specific variable. In the following explanation, we use the notation for a general function $f(\boldsymbol{x})$ to derive the proof. We wish to verify the following inequality for a class of functions $f(x_{j},\boldsymbol{x}_{C})$:
$\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{ice}(x_{j};\boldsymbol{x}_{C})\right]\geq I_{pdp}(x_{j}),$
or equivalently,
$\mathbb{E}_{\boldsymbol{x}_{C}}\left[\sqrt{\mathbb{V}_{x_{j}}[f(x_{j},\boldsymbol{x}_{C})]}\right]\geq\sqrt{\mathbb{V}_{x_{j}}\left[\mathbb{E}_{\boldsymbol{x}_{C}}[f(x_{j},\boldsymbol{x}_{C})]\right]}.$
Theorem 1.
If $f(x_{j},\boldsymbol{x}_{C})=g(x_{j})+h(\boldsymbol{x}_{C})$, where $g$ is a function of $x_{j}$ only and $h$ is a function of $\boldsymbol{x}_{C}$ only, then $\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{ice}(x_{j};\boldsymbol{x}_{C})\right]=I_{pdp}(x_{j})$.
Proof.
Let $f(x_{j},\boldsymbol{x}_{C})=g(x_{j})+h(\boldsymbol{x}_{C})$ where $g$ is a function of $x_{j}$ only and $h$ is a function of $\boldsymbol{x}_{C}$ only. First, note that
$\mathbb{E}_{\boldsymbol{x}_{C}}\left[\sqrt{\mathbb{V}_{x_{j}}[f(x_{j},\boldsymbol{x}_{C})]}\right]=\mathbb{E}_{\boldsymbol{x}_{C}}\left[\sqrt{\mathbb{V}_{x_{j}}[g(x_{j})+h(\boldsymbol{x}_{C})]}\right]=\mathbb{E}_{\boldsymbol{x}_{C}}\left[\sqrt{\mathbb{V}_{x_{j}}[g(x_{j})]}\right]=\sqrt{\mathbb{V}_{x_{j}}[g(x_{j})]}.$
Next, note that
$\sqrt{\mathbb{V}_{x_{j}}\left[\mathbb{E}_{\boldsymbol{x}_{C}}[f(x_{j},\boldsymbol{x}_{C})]\right]}=\sqrt{\mathbb{V}_{x_{j}}\left[\mathbb{E}_{\boldsymbol{x}_{C}}[g(x_{j})+h(\boldsymbol{x}_{C})]\right]}=\sqrt{\mathbb{V}_{x_{j}}\left[g(x_{j})+\mathbb{E}_{\boldsymbol{x}_{C}}[h(\boldsymbol{x}_{C})]\right]}=\sqrt{\mathbb{V}_{x_{j}}[g(x_{j})]}.$
Hence,
$\mathbb{E}_{\boldsymbol{x}_{C}}\left[\sqrt{\mathbb{V}_{x_{j}}[f(x_{j},\boldsymbol{x}_{C})]}\right]=\sqrt{\mathbb{V}_{x_{j}}\left[\mathbb{E}_{\boldsymbol{x}_{C}}[f(x_{j},\boldsymbol{x}_{C})]\right]}.$
Theorem 2.
If $f(x_{j},\boldsymbol{x}_{C})=g(x_{j})h(\boldsymbol{x}_{C})$, where $g$ is a function of $x_{j}$ only and $h$ is a function of $\boldsymbol{x}_{C}$ only, then $\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{ice}(x_{j};\boldsymbol{x}_{C})\right]\geq I_{pdp}(x_{j})$. Moreover, equality holds if $h$ is a nonnegative function.
Proof.
Let $f(x_{j},\boldsymbol{x}_{C})=g(x_{j})h(\boldsymbol{x}_{C})$ where $g$ is a function of $x_{j}$ only and $h$ is a function of $\boldsymbol{x}_{C}$ only. First, note that
[…]
Next, we have
$\begin{array}[]{rcl}\sqrt{\mathbb{V}_{x_{j}}\left[\mathbb{E}_{\boldsymbol{x}_{C}}[f(x_{j},\boldsymbol{x}_{C})]\right]}&=&\sqrt{\mathbb{V}_{x_{j}}\left[\mathbb{E}_{\boldsymbol{x}_{C}}[g(x_{j})h(\boldsymbol{x}_{C})]\right]}=\sqrt{\mathbb{V}_{x_{j}}\left[g(x_{j})\mathbb{E}_{\boldsymbol{x}_{C}}[h(\boldsymbol{x}_{C})]\right]}\\ \\ &=&\sqrt{(\mathbb{E}_{\boldsymbol{x}_{C}}[h(\boldsymbol{x}_{C})])^{2}\mathbb{V}_{x_{j}}[g(x_{j})]}=|\mathbb{E}_{\boldsymbol{x}_{C}}[h(\boldsymbol{x}_{C})]|\sqrt{\mathbb{V}_{x_{j}}[g(x_{j})]}.\end{array}$
By Jensen’s inequality, $\mathbb{E}_{\boldsymbol{x}_{C}}[|h(\boldsymbol{x}_{C})|]\geq|\mathbb{E}_{\boldsymbol{x}_{C}}[h(\boldsymbol{x}_{C})]|$, and so,
[…]
Moreover, if $h$ is a nonnegative function, then $\mathbb{E}_{\boldsymbol{x}_{C}}[h(\boldsymbol{x}_{C})]\geq 0$. In this case, $\mathbb{E}_{\boldsymbol{x}_{C}}[|h(\boldsymbol{x}_{C})|]=|\mathbb{E}_{\boldsymbol{x}_{C}}[h(\boldsymbol{x}_{C})]|$ and equality holds above. ∎
Before we state and prove the next theorem that covers a more general class of functions for which the above inequality holds, we first prove the following lemma.
Lemma 1.
If $A\in\mathbb{R}^{k\times k}$ is a symmetric and positive semidefinite matrix, then $\varphi(z)=\sqrt{z^{T}Az},\ z\in\mathbb{R}^{k}$ is a convex function. Moreover, if $Z$ is a $k$-dimensional random vector, then $E[\sqrt{Z^{T}AZ}]\geq\sqrt{E(Z)^{T}AE(Z)}$.
Proof.
Suppose $A$ is a symmetric positive semidefinite matrix. Then, $A=M^{T}M$ for some matrix $M\in\mathbb{R}^{k\times k}$ (e.g., $M$ may be obtained via spectral decomposition). Hence,
$\varphi(z)=\sqrt{z^{T}Az}=\sqrt{z^{T}M^{T}Mz}=\sqrt{(Mz)^{T}(Mz)}=\|Mz\|,$
where $\|\cdot\|$ is the 2-norm in $\mathbb{R}^{k}$. To prove the convexity of $\varphi(z)$, let $z_{1},z_{2}\in\mathbb{R}^{k}$ and let $0\leq\lambda\leq 1$. Note that
[…]
where the above inequality follows from the triangle inequality. This shows that
$\varphi(\lambda z_{1}+(1-\lambda)z_{2})\leq\lambda\varphi(z_{1})+(1-\lambda)\varphi(z_{2}),$
and so, $\varphi(z)$ is a convex function. Now, by Jensen’s Inequality, we have $E[\varphi(Z)]\geq\varphi(E[Z])$, or equivalently, $E[\sqrt{Z^{T}AZ}]\geq\sqrt{E(Z)^{T}AE(Z)}$. ∎
Theorem 3.
If $f(x_{j},\boldsymbol{x}_{C})=g_{1}(x_{j})h_{1}(\boldsymbol{x}_{C})+\ldots+g_{k}(x_{j})h_{k}(\boldsymbol{x}_{C})$, where $g_{1},\ldots,g_{k}$ are functions of $x_{j}$ only and $h_{1},\ldots,h_{k}$ are functions of $\boldsymbol{x}_{C}$ only, then $\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{ice}(x_{j};\boldsymbol{x}_{C})\right]\geq I_{pdp}(x_{j})$.
Proof.
Let $f(x_{j},\boldsymbol{x}_{C})=g_{1}(x_{j})h_{1}(\boldsymbol{x}_{C})+\ldots+g_{k}(x_{j})h_{k}(\boldsymbol{x}_{C})$, where $g_{1},\ldots,g_{k}$ are functions of $x_{j}$ only and $h_{1},\ldots,h_{k}$ are functions of $\boldsymbol{x}_{C}$ only. First, note that
[…]
Define the random vectors $G(x_{j})=[g_{1}(x_{j}),\ldots,g_{k}(x_{j})]^{T}$ and $H(\boldsymbol{x}_{C})=[h_{1}(\boldsymbol{x}_{C}),\ldots,h_{k}(\boldsymbol{x}_{C})]^{T}$. Now, note that
$\mathbb{V}_{x_{j}}[f(x_{j},\boldsymbol{x}_{C})]=\mathbb{V}_{x_{j}}[H(\boldsymbol{x}_{C})^{T}G(x_{j})]=H(\boldsymbol{x}_{C})^{T}\mbox{Cov}_{x_{j}}(G(x_{j}))H(\boldsymbol{x}_{C}),$
where
$\mbox{Cov}_{x_{j}}(G(x_{j}))=\begin{bmatrix}\mathbb{V}_{x_{j}}[g_{1}(x_{j})]&\mbox{Cov}_{x_{j}}(g_{1}(x_{j}),g_{2}(x_{j}))&\ldots&\mbox{Cov}_{x_{j}}(g_{1}(x_{j}),g_{k}(x_{j}))\\ \mbox{Cov}_{x_{j}}(g_{2}(x_{j}),g_{1}(x_{j}))&\mathbb{V}_{x_{j}}[g_{2}(x_{j})]&\ldots&\mbox{Cov}_{x_{j}}(g_{2}(x_{j}),g_{k}(x_{j}))\\ \vdots&\vdots&\ddots&\vdots\\ \mbox{Cov}_{x_{j}}(g_{k}(x_{j}),g_{1}(x_{j}))&\mbox{Cov}_{x_{j}}(g_{k}(x_{j}),g_{2}(x_{j}))&\ldots&\mathbb{V}_{x_{j}}[g_{k}(x_{j})]\\ \end{bmatrix}$
is the covariance matrix of $G(x_{j})$. Hence,
$\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{ice}(x_{j};\boldsymbol{x}_{C})\right]=\mathbb{E}_{\boldsymbol{x}_{C}}\left[\sqrt{\mathbb{V}_{x_{j}}[f(x_{j},\boldsymbol{x}_{C})]}\right]=\mathbb{E}_{\boldsymbol{x}_{C}}\left[\sqrt{H(\boldsymbol{x}_{C})^{T}\mbox{Cov}_{x_{j}}(G(x_{j}))H(\boldsymbol{x}_{C})}\right].$
Next, note that
[…]
Since $\mathbb{E}_{\boldsymbol{x}_{C}}[H(\boldsymbol{x}_{C})]=[\mathbb{E}_{\boldsymbol{x}_{C}}[h_{1}(\boldsymbol{x}_{C})],\ldots,\mathbb{E}_{\boldsymbol{x}_{C}}[h_{k}(\boldsymbol{x}_{C})]]^{T}$, we get
$\mathbb{V}_{x_{j}}\left[\mathbb{E}_{\boldsymbol{x}_{C}}[f(x_{j},\boldsymbol{x}_{C})]\right]=\mathbb{E}_{\boldsymbol{x}_{C}}[H(\boldsymbol{x}_{C})]^{T}\mbox{Cov}_{x_{j}}(G(x_{j}))\mathbb{E}_{\boldsymbol{x}_{C}}[H(\boldsymbol{x}_{C})],$
and so
$I_{pdp}(x_{j})=\sqrt{\mathbb{V}_{x_{j}}\left[\mathbb{E}_{\boldsymbol{x}_{C}}[f(x_{j},\boldsymbol{x}_{C})]\right]}=\sqrt{\mathbb{E}_{\boldsymbol{x}_{C}}[H(\boldsymbol{x}_{C})]^{T}\mbox{Cov}_{x_{j}}(G(x_{j}))\mathbb{E}_{\boldsymbol{x}_{C}}[H(\boldsymbol{x}_{C})]}.$
Now, using the previous lemma with $Z=H(\boldsymbol{x}_{C})$ and with $A=\mbox{Cov}_{x_{j}}(G(x_{j}))$, we obtain
[…]
Corollary 1.
If $f$ is a polynomial in $x_{1},\ldots,x_{m}$, then $\mathbb{E}_{\boldsymbol{x}_{C}}\left[I_{ice}(x_{j};\boldsymbol{x}_{C})\right]\geq I_{pdp}(x_{j})$ for any $j=1,\ldots,m$.
Proof.
If $f$ is a polynomial in $x_{1},\ldots,x_{m}$, then for any $j=1,\ldots,m$, $f$ has the form $f(x_{j},\boldsymbol{x}_{C})=g_{1}(x_{j})h_{1}(\boldsymbol{x}_{C})+\ldots+g_{k}(x_{j})h_{k}(\boldsymbol{x}_{C})$, where $g_{1},\ldots,g_{k}$ are functions of $x_{j}$ only and $h_{1},\ldots,h_{k}$ are functions of $\boldsymbol{x}_{C}$ only. The result follows from the previous theorem. ∎
The above corollary shows that the ICE-PDP inequality holds for all multivariate polynomials. However, it holds for a much larger class of functions that has some combination of additive and multiplicative separability, such as
$f(x_{1},x_{2})=x_{1}^{2}\sin(\pi x_{2})+e^{x_{1}}{x_{2}}^{3}+\cos(\pi x_{1})\sqrt{x_{2}}.$
We conjecture that the inequality also holds for more general functions, though an investigation of this claim will be left for future work.